OpenStudy (anonymous):

Problem for you: Find the period of the function in the simplest way:

OpenStudy (anonymous):

\[f(x+3) + f(x-3) = \sqrt{3} f(x)\]

OpenStudy (anonymous):

Yeah a bit nasty

OpenStudy (anonymous):

Haha. It has an elegant solution, like always. Best of luck.

OpenStudy (anonymous):

Well I think you will excuse me (and may be even ask one of your pals to mdl me, since I want to leave the full soln to others) \[ (e^{-i3\nu} + e^{i3\nu} ) \cal{F} (\nu) = \sqrt{3} \cal{F} (\nu) \]

OpenStudy (anonymous):

\[ \frac{2}{\sqrt3}Cos(3\nu)\cal{F} (\nu) = \cal{F} (\nu)\]

OpenStudy (anonymous):

(1)I do not understand how you made that conversion. (2) This is elementary high school maths. Keep it simple.

OpenStudy (anonymous):

let x=x+3\[f(x+6)+f(x)=\sqrt{3} \ f(x-3)\]let x=x-3\[f(x)+f(x-6)=\sqrt{3} \ f(x+3)\]add them\[f(x+6)+f(x-6)+2f(x)=\sqrt{3} \ (f(x-3)+f(x+3))=\sqrt{3} \ (\sqrt{3} \ f(x))=3f(x)\]

OpenStudy (anonymous):

so\[f(x+6)+f(x-6)=f(x)\]

OpenStudy (anonymous):

let x=x+3\[f(x+9)+f(x-3)=f(x+3)\]let x=x-3\[f(x+3)+f(x-9)=f(x-3)\]add them\[f(x+9)+f(x-9)=0\]period is 18

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

emm...maybe ??? check my work...

OpenStudy (anonymous):

Noo. Its all correct. Just the conclusion doesnt seem right.

OpenStudy (anonymous):

wait a min

OpenStudy (anonymous):

Period is T when: \[f(x) = f(x +T)\]

OpenStudy (anonymous):

\[f(x+18)=-f(x)\]let x=x+9\[f(x+27)=-f(x+9)=f(x-9)\]period is 36

OpenStudy (anonymous):

i think we are done

OpenStudy (anonymous):

Precisely. Well done!!

OpenStudy (anonymous):

@Mikael See. No Fourier's involved. :)

OpenStudy (anonymous):

Still - please prove or DISprove the equation above. I would be mcuh obliged in either case. BTW u all should be intrigued by it - since it says f =0 identically @mukushla @siddhantsharan

OpenStudy (anonymous):

Yes, but as is well known, the undeclared but common interpretation of functional equations like that is that they hold FOR ALL real x/domain of definition